OpenStudy (anonymous):

Solve the equation. x^2 + 18x + 81 = 25 A. 14, 4 B. –4, –14 C. 14, –14 D. –4, 4 my answer is B

OpenStudy (anonymous):

can some one tell if im correct

OpenStudy (helder_edwin):

yes u r.

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

\[ (x+9)^2 = 25 \]Do you see how I got that? Inspection! From here, break up into two equations: x + 9 = 25 and x + 9 = -25. Boom!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem. Thanks for the medal! :D